OpenStudy (anonymous):

-4x+7y=-14 what is the slope intercept form

OpenStudy (anonymous):

All you have to do is get y by itself. You need everything in y=mx+b form.

OpenStudy (anonymous):

I tried that, but the answer does not look correct. I cam up with y=-4/7x+2

OpenStudy (anonymous):

-4x+7y=-14 +4x +4x ________________________ 7y= 4x-14 Now get y by itself by dividing 7 on both sides. y=\[\frac{ 4 }{ 7 }x-\frac{ 14? }{ 7 }\] NOW SIMPLIFY! :) HOPE THIS HELPED!

OpenStudy (anonymous):

\[\frac{ -14 }{ 7 }\neq positive 2\]